# Explainability-Based Token Replacement on LLM-Generated Text

### v1

@@ Abstract.
Results: Our findings show that these token replacement approaches can significantly diminish a single classifier’s ability to detect AIGT. However, our ensemble classifier maintains strong performance across multiple languages and domains, showing that a multi-model approach can mitigate the impact of token-level manipulations.
@@ 1. Introduction
Recent advancements in LLMs, such as the GPT series, have enabled the creation of highly fluent text that often resembles human writing (Brown et al., 2020). Still, even the best models often show subtle signs of being artificial, like repeating phrases, using a limited range of words, having no typos, or showing a consistent writing style. These clues can help readers or automated tools detect that the text was written by AI (Sadasivan et al., 2023). Being able to spot AIGT is important in various domains, such as in for protecting academic honesty, fighting misinformation, and keeping public trust. On the other hand, there are good reasons to make AI text less detectable, like tools that help people write in their own style (Yeh et al., 2024), or AI summaries that aim to be especially clear and well-structured (Nacke et al., 2024).
This conflict between helpful and harmful uses of LLMs creates a challenge: how to keep their benefits while preventing misuse. One proposed solution is watermarking, which adds hidden patterns to generated text to make it more transparent (Kirchenbauer et al., 2023b). However, recent research shows that these watermarks can be weakened or removed by changing the text, like paraphrasing (Li et al., 2023). Against this backdrop, our study asks: Can explanation methods be used to both hide and detect AI-generated text? We contribute in two ways. First, we develop an ensemble‑based detector that remains accurate across languages (English and Dutch) and domains (news, reviews, Twitter). Second, we show how two explanation tools, SHapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017) and Local Interpretable Model-agnostic Explanations (LIME) (Ribeiro et al., 2016), can be turned against a detector: we identify the most influential tokens and replace them via four targeted strategies (semantically similar human‑preferred words, part‑of‑speech‑constrained synonyms, GPT‑based substitutions, and GPT substitutions enriched with genre context). We evaluate both detectability (accuracy, F1) and textual fidelity (BLEU, ROUGE), thereby quantifying the trade‑off between hiding AI signals and preserving meaning. Our results show that token‑level manipulation can fool many single‑model detectors, yet a robust ensemble still performs strongly, underscoring the need to combine explainability, watermarking, and policy safeguards to manage the evolving landscape of AIGT.
@@ 2. Related Work
Kirchenbauer et al. (2023b) introduced an important red–green list watermark. In this method, a secret hash divides the vocabulary at each token position into a preferred “green” group and a less preferred “red” group. The model then gives higher chances to the green tokens, so the generated text contains a hidden cryptographic signature. Light editing often keeps these watermarks, but bigger changes, like paraphrasing or replacing words with synonyms, can remove them (Kirchenbauer et al., 2023a; Li et al., 2023). To make watermarks harder to notice, distortion-free methods hide information in the token-sampling order instead of the probability distribution. However, these can also become weak if someone rewrites the text in a tricky way (Christ et al., 2024). Other approaches, called post-hoc schemes, add watermarks later by changing syntax templates or choosing context-aware synonyms (Yang et al., 2023a), but they also have trouble surviving strong text changes.
@@
Another important detection method takes advantage of the fact that language models (LMs) often choose very predictable words. Gehrmann et al. (2019) created a token-rank histogram tool to help people find words that are “too likely” in a sentence. Zero-shot methods like DetectGPT look at the “curvature” of local log-probabilities; if the curvature is negative, it suggests the text was written by a machine (Mitchell et al., 2023). Fast-DetectGPT makes this process faster and more accurate for text from ChatGPT or GPT-4 (Bao et al., 2023). Other detection methods use features such as normalized perturbed rank (Su et al., 2023) and the amount of unused information in token probability distributions (Zawistowski, 2024). These “white-box” methods work well when they can directly access the model’s probabilities, but they become less effective without this access. To solve this problem, “DNA” regeneration methods, like DNA-GPT (Yang et al., 2023b) and GPT Paternity Test (Yu et al., 2023), do not need internal access; instead, they query a black-box model again and compare the new outputs for similarity.
@@
Stylometric approaches instead focus on word choice, sentence structure, and how the text is organized and connected. Fröhling and Zubiaga (2021) grouped 35 linguistic features into four categories, low diversity, repetition, incoherence, and lack of purpose, and trained SVM and random-forest models. More advanced work introduces entity-coherence graphs (Liu et al., 2023) or journalism-specific style metrics (Kumarage et al., 2023), which bolster detection on brief social-media texts. Though interpretable, these handcrafted methods may be narrower in scope and less transferable across domains.
@@
A different direction uses fine-tuned neural networks to classify AI-generated text. OpenAI’s initial GPT-detector trained RoBERTa on mixed GPT-2 corpora, achieving roughly 95% accuracy on in-domain data (Solaiman et al., 2019). More recent detectors employ Longformer (Beltagy et al., 2020) or ELECTRA to handle longer contexts and detect texts from multiple generators (Clark et al., 2020). However, these systems often struggle with out-of-distribution domains. Recent “cross-model” or contrastive ensembles, such as Ghostbuster (Verma et al., 2023) and DeTeCtive (Guo et al., 2024), compare probability distributions from multiple open-source LMs to improve robustness.
@@
Ensemble methods combine diverse signals to mitigate the weaknesses of individual detectors. Simple voting stacks of transformer-based classifiers outperform single models in multilingual tasks (Nguyen et al., 2023). For example, recent work introduces ensemble-based approaches for detecting AI-generated text in educational content, achieving a favorable false-positive trade-off in academic scenarios (Najjar et al., 2025). Our work follows this line by designing an explanation-oriented ensemble meant to withstand token-level adversarial edits.
@@
Paraphrasing is one of the best ways to trick AI detectors. For example, Krishna et al. (2023) showed that changing the wording can make DetectGPT miss almost all AI-written text. Rewriting the text many times makes it even harder to detect (Sadasivan et al., 2023). Some new methods, like RADAR (Hu et al., 2023) and OUTFOX (Koike et al., 2024), are better at catching these changes because they are trained on paraphrased examples, but they still only recover some accuracy. Even small mistakes, like typos, can also confuse detectors (Huang et al., 2024). More advanced methods can handle some edits, but none can stop all tricks or rewrites.
@@
Paraphrasing is seen as the strongest way to avoid AI text detectors: Krishna et al. (2023) found that rewriting text can reduce DetectGPT’s ability to catch AI-generated text from 70% to less than 5%. Rewriting the text several times in a row (called “self-rewrites”) makes it even harder to detect, while still keeping the original meaning (Sadasivan et al., 2023). Defenses such as RADAR (Hu et al., 2023) and OUTFOX (Koike et al., 2024) become stronger by training on paraphrased texts, and can recover 10–20 F1 points. Recent studies also show that even simple changes at the letter level—like typos or swapping similar-looking letters—can confuse detectors and make them much less accurate (Huang et al., 2024). Although using features like ensemble cross-probability or advanced watermarking can resist some small changes, no method is completely safe from strong rewriting attacks.
@@
Explainability has become a cornerstone of modern NLP deployments (Mohammadi et al., 2025). XAI tools like SHAP and LIME show which words or tokens have the biggest effect on a detector’s decisions. For example,  Mitrovi’c et al. (2023) used LIME to highlight important words and help people review short ChatGPT answers. But attackers can use this information in the opposite way, by changing or removing the most important words. Zhou et al. (2024) showed that using SHAP-based synonyms for key words can lower detector accuracy by 40–60%. In our work, we carefully study this double-sided use of explanations, testing four types of word substitutions in two languages and different types of texts.
@@
Detectors often make mistakes with texts written by people using English as a second language. Liang et al. (2023) found that seven popular GPT detectors wrongly marked more than 61% of TOEFL essays by non-native English speakers as AI-generated, even though they were very accurate with essays by native speakers. Because of this bias, some teachers are moving away from strict rules and are now encouraging students to use AI tools as support, especially for those with different language backgrounds. Since LLMs can quickly create lots of convincing fake news, it is very hard for people alone to spot which content is real (Zellers et al., 2019). To improve detection in these situations, it helps to use a mix of classifiers and to add information about where the text comes from (Fraser et al., 2025).
@@
Recent use of AI chatbots for mental health support shows that people often think these systems are caring and human-like, which leads to real and meaningful conversations. But this can be risky, especially when users do not realize they are talking to an AI.  Song et al. (2024) found that people in severe distress sometimes seek help from chatbots and even see them as human. In the same way, Siddals et al. (2024) stressed the need for stronger safety rules and clearer information in AI mental health tools. These studies show how important it is to have good detection systems and to make it clear when someone is interacting with an AI, so that chatbots are used ethically in sensitive cases. Research also shows that no single method is perfect; reliable AI-generated text detection needs to use many signals, be tested against tricky cases, and give clear reasons for its decisions. These are the main goals of our work.
@@ 3. Dataset
In this section, we detail the CLIN33 dataset (Fivez et al., 2024) used as our primary corpus and describe how it was augmented to expand its linguistic variability. We also present the portion of the AuTexTification dataset [^1] https://zenodo.org/records/10732813 used to strengthen our ensemble models. The CLIN33 shared task corpus comprises human-written and AIGTs in English and Dutch, including three domains: news, tweets, and reviews. Each domain contains 200 samples from human authors and 200 samples from LLMs (GPT-4 and Vicuña-13B), producing a total of 1,200 texts in English and 1,200 in Dutch. The human-written news samples come from well-known media outlets; the Dutch social media posts primarily discuss policy or health topics, whereas the English ones often revolve around social or sports events. Reviews typically cover literature in Dutch and various consumer goods in English.
Although CLIN33 offers balanced classes across two languages and three genres, its size remains relatively small for training large-scale models. To address this limitation and diversify the textual patterns, we created an augmented version of the dataset. We started with the original texts and applied a series of transformations to each sample. These transformations included synonym substitution, random word swaps, random insertions, targeted deletions, variations in spelling, and back-translation [^2] Translating from English to Dutch and then back to English. Each text thus produced several new variants that preserved the key semantics but introduced new surface forms. We ultimately obtained 9,720 total samples for training, split evenly between human and AI-generated content (4,860 each) and distributed equally across the three domains and the two languages. The final augmented corpus enables more robust feature learning.
@@
Augmentations None Yes (9)
As an additional resource for pre-training and fine-tuning, we employed a subset of the AuTexTification dataset [^3] https://sites.google.com/view/autextification, introduced by Sarvazyan et al. (2023). While the complete dataset includes English and Spanish texts in five distinct domains, we selectively used the English-only portion from tweets, legal, and wiki categories. This subset contains 33,845 texts designated for binary AI-detection (Task 1), split nearly evenly between human-written (50.36%) and AI-generated (49.64%) documents. The human content is drawn from sources such as MultiEURLEX, XSUM, Amazon Reviews, TSATC, and WikiLingua, while the AI-generated portion is produced by BLOOM (1B7, 3B, 7B1) and GPT-3 (babbage, curie, text-davinci-003). Because of its class balance and domain variety, this subset of AuTexTification serves as a valuable source of pre-training examples for our ensemble. Table 1 summarizes the essential attributes of the two datasets used to train and evaluate our methodology. The combined setup draws from the 33,845 samples in AuTexTification (English only) and the 9,720 augmented CLIN33 samples (both English and Dutch). We use AuTexTification Dataset for initial pre-training, 90% of the CLIN data for the Augmenting and training and 10% of unchanged CLIN data for evaluating.
4. Methodology
We used an approach that integrates several models for text classification to address the AIGT detection task. First, we converted all text to lowercase to handle tokens in a consistent way. We then removed non-informative elements such as URLs, punctuation, and special characters, since these tokens often do not carry semantic meaning. After this cleaning, we split the text into tokens and applied lemmatization to map words to their base forms, which lowered data complexity and improved the model’s ability to recognize patterns.
Baselines Models.
We built an XGBoost classifier by first transforming the text into numerical features with a TF-IDF vectorizer set to a maximum of 5000 features. The XGBoost model was configured with use_label_encoder=False and eval_metric=’logloss’. Training used the TF-IDF representations from the training set, and final predictions on the test set were evaluated using precision, recall, F1, and accuracy. Also, We fine-tuned BERT-base (bert-base-uncased), DistilBERT (distilbert-base-uncased), and XLM-RoBERTa (xlm-roberta-base). The best checkpoint was chosen based on the highest F1 score on a validation subset. We then evaluated the final models on the test data.
@@ 4. Methodology
[…]
@@
[…]
@@
[…]

### v2

@@ Abstract.
Results: Our findings show that these token replacement approaches can substantially diminish a single classifier’s ability to detect AIGT. Human evaluators achieved only 47.0% overall detection accuracy, showing that our strategies deceive both machines and humans. However, our ensemble classifier maintains strong performance across multiple languages and domains, showing that a multi-model approach can mitigate the impact of token-level manipulations.
@@ 1. Introduction
Recent advancements in LLMs, such as the GPT series, have enabled the creation of highly fluent text that often resembles human writing (brown2020language). Still, even the best models often show subtle signs of being artificial, like repeating phrases, using a limited range of words, having no typos, or showing a consistent writing style. These clues can help readers or automated tools detect that the text was written by AI (sadasivan2023can). Being able to spot AIGT is important in various domains, such as in for protecting academic honesty, fighting misinformation, and keeping public trust. On the other hand, there are good reasons to make AI text less detectable, like tools that help people write in their own style (Yeh2024), or AI summaries that aim to be especially clear and well-structured (nacke2024ai).
This conflict between helpful and harmful uses of LLMs creates a challenge: how to keep their benefits while preventing misuse. One proposed solution is watermarking, which adds hidden patterns to generated text to make it more transparent (kirchenbauer2023watermarking). However, recent research shows that these watermarks can be weakened or removed by changing the text, like paraphrasing (li2023towards). Against this backdrop, our study asks: Can explanation methods be used to both hide and detect AI-generated text? We contribute in two ways. First, we develop an ensemble‑based detector that remains accurate across languages (English and Dutch) and domains (news, reviews, Twitter). Second, we show how two explanation tools, SHapley Additive exPlanations (SHAP) (lundberg2017unified) and Local Interpretable Model-agnostic Explanations (LIME) (ribeiro2016should), can be turned against a detector: we identify the most influential tokens and replace them via four targeted strategies (semantically similar human‑preferred words, part‑of‑speech‑constrained synonyms, GPT‑based substitutions, and GPT substitutions enriched with genre context). We evaluate both detectability (accuracy, F1) and textual fidelity (BLEU, ROUGE), thereby quantifying the trade‑off between hiding AI signals and preserving meaning. Our results show that token‑level manipulation can fool many single‑model detectors, yet a robust ensemble still performs strongly, underscoring the need to combine explainability, watermarking, and policy safeguards to manage the evolving landscape of AIGT.
@@ 2. Related Work
kirchenbauer2023watermarking introduced an important red–green list watermark. In this method, a secret hash divides the vocabulary at each token position into a preferred “green” group and a less preferred “red” group. The model then gives higher chances to the green tokens, so the generated text contains a hidden cryptographic signature. Light editing often keeps these watermarks, but bigger changes, like paraphrasing or replacing words with synonyms, can remove them (kirchenbauer2024reliability; li2023towards). To make watermarks harder to notice, distortion-free methods hide information in the token-sampling order instead of the probability distribution. However, these can also become weak if someone rewrites the text in a tricky way (christ2024undetectable). Other approaches, called post-hoc schemes, add watermarks later by changing syntax templates or choosing context-aware synonyms (yang2023watermarking), but they also have trouble surviving strong text changes.
@@
Another important detection method takes advantage of the fact that language models (LMs) often choose very predictable words. gehrmann2019gltr created a token-rank histogram tool to help people find words that are “too likely” in a sentence. Zero-shot methods like DetectGPT look at the “curvature” of local log-probabilities; if the curvature is negative, it suggests the text was written by a machine (mitchell2023detectgpt). Fast-DetectGPT makes this process faster and more accurate for text from ChatGPT or GPT-4 (bao2023fast). Other detection methods use features such as normalized perturbed rank (su2023detectllm) and the amount of unused information in token probability distributions (zawistowski2024unused). These “white-box” methods work well when they can directly access the model’s probabilities, but they become less effective without this access. To solve this problem, “DNA” regeneration methods, like DNA-GPT (yang2023dna) and GPT Paternity Test (yu2023gpt), do not need internal access; instead, they query a black-box model again and compare the new outputs for similarity.
@@
Stylometric approaches instead focus on word choice, sentence structure, and how the text is organized and connected. frohling2021feature grouped 35 linguistic features into four categories, low diversity, repetition, incoherence, and lack of purpose, and trained SVM and random-forest models. More advanced work introduces entity-coherence graphs (liu2023coco) or journalism-specific style metrics (kumarage2023j), which bolster detection on brief social-media texts. Though interpretable, these handcrafted methods may be narrower in scope and less transferable across domains.
@@
A different direction uses fine-tuned neural networks to classify AI-generated text. OpenAI’s initial GPT-detector trained RoBERTa on mixed GPT-2 corpora, achieving roughly 95% accuracy on in-domain data (solaiman2019release). More recent detectors employ Longformer (beltagy2020longformer) or ELECTRA to handle longer contexts and detect texts from multiple generators (clark2020electra). However, these systems often struggle with out-of-distribution domains. Recent “cross-model” or contrastive ensembles, such as Ghostbuster (verma2023ghostbuster) and DeTeCtive (guo2024detective), compare probability distributions from multiple open-source LMs to improve robustness.
@@
Ensemble methods combine diverse signals to mitigate the weaknesses of individual detectors. Simple voting stacks of transformer-based classifiers outperform single models in multilingual tasks (nguyen2023stacking). For example, recent work introduces ensemble-based approaches for detecting AI-generated text in educational content, achieving a favorable false-positive trade-off in academic scenarios (najjar2025detecting). Our work follows this line by designing an explanation-oriented ensemble meant to withstand token-level adversarial edits.
@@
Paraphrasing is one of the best ways to trick AI detectors. For example, krishna2023paraphrasing showed that changing the wording can make DetectGPT miss almost all AI-written text. Rewriting the text many times makes it even harder to detect (sadasivan2023can). Some new methods, like RADAR (hu2023radar) and OUTFOX (koike2024outfox), are better at catching these changes because they are trained on paraphrased examples, but they still only recover some accuracy. Even small mistakes, like typos, can also confuse detectors (huang2024ai). More advanced methods can handle some edits, but none can stop all tricks or rewrites.
@@
Paraphrasing is seen as the strongest way to avoid AI text detectors: krishna2023paraphrasing found that rewriting text can reduce DetectGPT’s ability to catch AI-generated text from 70% to less than 5%. Rewriting the text several times in a row (called ”self-rewrites”) makes it even harder to detect, while still keeping the original meaning (sadasivan2023can). Defenses such as RADAR (hu2023radar) and OUTFOX (koike2024outfox) become stronger by training on paraphrased texts, and can recover 10–20 F1 points. Recent studies also show that even simple changes at the letter level—like typos or swapping similar-looking letters—can confuse detectors and make them much less accurate (huang2024ai). Beyond paraphrase-based evasion, reinforcement learning has emerged as a new paradigm for generating detector-resistant text. AuthorMist (david2025authormist) employs RL to iteratively refine text modifications that maximize evasion success while preserving semantic content, demonstrating that adaptive adversaries can systematically exploit detector weaknesses. Although using features like ensemble cross-probability or advanced watermarking can resist some small changes, no method is completely safe from strong rewriting attacks.
@@
Explainability has become a cornerstone of modern NLP deployments (mohammadi2025explainability). XAI tools like SHAP and LIME show which words or tokens have the biggest effect on a detector’s decisions. For example,  mitrovic2023chatgpt used LIME to highlight important words and help people review short ChatGPT answers. But attackers can use this information in the opposite way, by changing or removing the most important words. zhou2024humanizing showed that using SHAP-based synonyms for key words can lower detector accuracy by 40–60%. In our work, we carefully study this double-sided use of explanations, testing four types of word substitutions in two languages and different types of texts.
Beyond SHAP and LIME, alternative explainability approaches have been explored for NLP tasks. SyntaxShap (amara2024syntaxshap) extends SHAP by incorporating syntactic structure, assigning importance scores to phrase-level constituents rather than individual tokens, which can provide more linguistically meaningful explanations for text generation tasks. Similarly, Integrated Gradients has been applied to detect adversarial attacks on text classifiers (moraliyage2025explainable), offering gradient-based attribution that traces predictions back through neural network layers. We chose SHAP and LIME for our study because they are model-agnostic and widely applicable to any classifier architecture, including our ensemble of transformers and XGBoost models. This flexibility allows practitioners to apply our token replacement framework regardless of the underlying detection model.
A natural question arises regarding the relationship between detection-based approaches and watermarking schemes. While watermarking proactively embeds identifiable patterns during text generation (kirchenbauer2023watermarking), detection methods operate post-hoc on any text regardless of its origin. These approaches are complementary: watermarking provides provenance verification for known LLM outputs, while robust detection remains necessary for texts generated without watermarks or by adversaries who actively remove them (li2023towards). Our XAI-based approach could potentially enhance watermarking schemes by identifying which tokens carry the strongest detection signal, informing where watermarks might be most resilient to modification. This synergy between explainability, detection, and watermarking represents a promising direction for comprehensive AI text governance.
@@
Detectors often make mistakes with texts written by people using English as a second language. liang2023gpt found that seven popular GPT detectors wrongly marked more than 61% of TOEFL essays by non-native English speakers as AI-generated, even though they were very accurate with essays by native speakers. Because of this bias, some teachers are moving away from strict rules and are now encouraging students to use AI tools as support, especially for those with different language backgrounds. Since LLMs can quickly create lots of convincing fake news, it is very hard for people alone to spot which content is real (zellers2019defending). To improve detection in these situations, it helps to use a mix of classifiers and to add information about where the text comes from (fraser2025detecting).
@@
Recent use of AI chatbots for mental health support shows that people often think these systems are caring and human-like, which leads to real and meaningful conversations. But this can be risky, especially when users do not realize they are talking to an AI.  song2024typing found that people in severe distress sometimes seek help from chatbots and even see them as human. In the same way, siddals2024happened stressed the need for stronger safety rules and clearer information in AI mental health tools. These studies show how important it is to have good detection systems and to make it clear when someone is interacting with an AI, so that chatbots are used ethically in sensitive cases. Research also shows that no single method is perfect; reliable AI-generated text detection needs to use many signals, be tested against tricky cases, and give clear reasons for its decisions. These are the main goals of our work.
@@ 3. Dataset
In this section, we detail the CLIN33 dataset (fivez2024clin33) used as our primary corpus and describe how it was augmented to expand its linguistic variability. We also present the portion of the AuTexTification dataset [^1] https://zenodo.org/records/10732813 used to strengthen our ensemble models. The CLIN33 shared task corpus comprises human-written and AIGTs in English and Dutch, including three domains: news, tweets, and reviews. Each domain contains 200 samples from human authors and 200 samples from LLMs (GPT-4 and Vicuña-13B), producing a total of 1,200 texts in English and 1,200 in Dutch. The human-written news samples come from well-known media outlets; the Dutch social media posts primarily discuss policy or health topics, whereas the English ones often revolve around social or sports events. Reviews typically cover literature in Dutch and various consumer goods in English.
@@
[…]
Although CLIN33 offers balanced classes across two languages and three genres, its size remains relatively small for training large-scale models. To address this limitation and diversify the textual patterns, we created an augmented version of the dataset. We started with the original texts and applied a series of transformations to each sample. These transformations included synonym substitution, random word swaps, random insertions, targeted deletions, variations in spelling, and back-translation [^2] Translating from English to Dutch and then back to English. Each text thus produced several new variants that preserved the key semantics but introduced new surface forms. We generated an augmented pool of 9,720 instances by applying surface-level transformations (e.g., synonym substitution, swaps/insertions/deletions, spelling variants, and back-translation) to CLIN33 texts. From the 2,160 training texts (90% of CLIN33), we applied transformations to create an average of 3.5 augmented variants per original text. Including the original texts, this yielded 9,720 total training instances (4.5 instances per source text on average). To avoid test leakage, augmentation was applied only to the training split (described below), while the held-out test set remained non-augmented.
As an additional resource for pre-training and fine-tuning, we employed a subset of the AuTexTification dataset [^3] https://sites.google.com/view/autextification, introduced by sarvazyan2023overview. While the complete dataset includes English and Spanish texts in five distinct domains, we selectively used the English-only portion from tweets, legal, and wiki categories. This subset contains 33,845 texts designated for binary AI-detection (Task 1), split nearly evenly between human-written (50.36%) and AI-generated (49.64%) documents. The human content is drawn from sources such as MultiEURLEX, XSUM, Amazon Reviews, TSATC, and WikiLingua, while the AI-generated portion is produced by BLOOM (1B7, 3B, 7B1) and GPT-3 (babbage, curie, text-davinci-003). Because of its class balance and domain variety, this subset of AuTexTification serves as a valuable source of pre-training examples for our ensemble. Table 1 summarizes the essential attributes of the two datasets used to train and evaluate our methodology. The combined setup draws from the 33,845 samples in AuTexTification (English only) and the 9,720 augmented CLIN33 samples (both English and Dutch). We use AuTexTification Dataset for initial pre-training. We use 90% of the original CLIN33 texts to construct training data (including augmentation), and reserve the remaining 10% of original CLIN33 as an unchanged held-out test set.
We first created a stratified 90/10 split on the original (non-augmented) CLIN33 texts, preserving balance across language $\times$ domain $\times$ label. The held-out test set contains 240 original texts (120 human, 120 AI): 120 English and 120 Dutch, and 40 texts per domain per language (20 human, 20 AI). The remaining 90% of the original CLIN33 texts were used for training/validation, and only that portion was augmented.
@@
2:  $\text{bert\_out}\leftarrow\text{BERTModel}_{\text{frozen}}(x)$
3:  $\text{xlmr\_out}\leftarrow\text{XLMRobertaModel}_{\text{frozen}}(x)$
4:  $\text{distil\_out}\leftarrow\text{DistilBERTModel}_{\text{frozen}}(x)$
5:  return $(\text{bert\_out},\text{xlmr\_out},\text{distil\_out})$
@@
8:  $\text{bert\_out\_fresh}\leftarrow\text{BERTModel}_{\text{fresh}}(x)$
9:  $\text{xlmr\_out\_fresh}\leftarrow\text{XLMRobertaModel}_{\text{fresh}}(x)$
10:  $\text{distil\_out\_fresh}\leftarrow\text{DistilBERTModel}_{\text{fresh}}(x)$
11:  return $(\text{bert\_out\_fresh},\text{xlmr\_out\_fresh},\text{distil\_out\_fresh})$
@@
14:  $(\text{froz\_bert},\text{froz\_xlmr},\text{froz\_distil})\leftarrow\textsc{PretrainEnsemble}(x)$
15:  $(\text{fresh\_bert},\text{fresh\_xlmr},\text{fresh\_distil})\leftarrow\textsc{FineTuneEnsemble}(x)$
16:  $\text{concat\_outs}\leftarrow\text{concatenate}[\text{froz\_bert},\text{froz\_xlmr},\text{froz\_distil},$
17:    $\text{fresh\_bert},\text{fresh\_xlmr},\text{fresh\_distil}]$
18:  $\text{dense\_out}\leftarrow\text{DenseLayer}(\text{concat\_outs})$
19:  $\hat{y}\leftarrow\sigma(\text{dense\_out})$
@@
4. Methodology
We used an approach that integrates several models for text classification to address the AIGT detection task. For XGBoost we apply lowercase/URL removal/punctuation removal/lemmatization. For transformer models we use the native tokenizer with no extra preprocessing. We fine-tune bert-base-multilingual-uncased, distilbert-base-multilingual-cased, and xlm-roberta-base for both English and Dutch.
Baselines Models.
We built an XGBoost classifier by first transforming the text into numerical features with a TF-IDF vectorizer set to a maximum of 5000 features. The XGBoost model was configured with use_label_encoder=False and eval_metric=’logloss’. Training used the TF-IDF representations from the training set, and final predictions on the test set were evaluated using precision, recall, F1, and accuracy. Also, We fine-tuned BERT-base (bert-base-multilingual-uncased), DistilBERT (distilbert-base-multilingual-cased), and XLM-RoBERTa (xlm-roberta-base). The best checkpoint was chosen based on the highest F1 score on a validation subset. We then evaluated the final models on the test data.